Ln x - Ln(x-9)=Ln 9
Answer by oscargut(686)

Ln x - Ln(x-9)=Ln 9
then
ln(x/(x-9)=Ln9
x/(x-9)=9
then
x=9(x-9)=9x-81
8x-81=0
x=81/8